# Benefits of Mutual Coupling in Dynamic Metasurface Antennas for Optimizing Wireless Communications – Theory and Experimental Validation

### v1

@@ II-C Insights into Mutual Coupling
$\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})=\left[\left(\mathbf{W}(\mathbf{c})\right)^{-1}\right]_{\mathcal{MM}}=\left(\tilde{\mathbf{W}}-\mathrm{diag}(\mathbf{c})\right)^{-1},$ (8b)
@@
$\begin{split}\left[\left(\mathbf{W}(\mathbf{c})\right)^{-1}\right]_{\mathcal{MF}}=\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})&\mathbf{W}^{0}_{\mathcal{MF}}\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}\\ =\left[\tilde{\mathbf{W}}^{-1}-\tilde{\mathbf{W}}^{-1}\left(\mathbf{I}_{N_{\mathrm{M}}}+\tilde{\mathbf{W}}\mathrm{diag}(\mathbf{c})\right)^{-1}\right]&\mathbf{W}^{0}_{\mathcal{MF}}\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}.\end{split}$ (9)
Substituting Eq. (9) into Eq. (7), we obtain
@@
$\begin{split}e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})=\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\!\left(\tilde{\mathbf{W}}\!-\!\mathrm{diag}(\mathbf{c})\right)^{-1}&\mathbf{W}^{0}_{\mathcal{MF}}\!\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}\!\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}}\\ =\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\!\tilde{\mathbf{W}}^{-1}&\mathbf{W}^{0}_{\mathcal{MF}}\!\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}\!\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}}\\ -\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\!\tilde{\mathbf{W}}^{-1}\!\left(\mathbf{I}_{N_{\mathrm{M}}}\!+\!\tilde{\mathbf{W}}\mathrm{diag}(\mathbf{c})\right)^{-1}&\mathbf{W}^{0}_{\mathcal{MF}}\!\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}\!\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}}.\end{split}$ (10)
@@
The first [second] term on the right hand side in Eq. (10) represents contributions to the radiated field that are independent [dependent] of the DMA configuration $\mathbf{c}$. The dependence on $\mathbf{c}$ is seen to generally be non-linear due to the involved matrix inversion. By inspection of Eq. (8b), we observe that the off-diagonal elements of $\tilde{\mathbf{W}}$ mediate the coupling between the different entries of $\mathbf{c}$. This mathematical observation precisely matches the physical interpretation. The off-diagonal elements of the first term on the right hand side in Eq. (8a) contain the background Green’s functions between the meta-atoms which capture the coupling between them that is mediated via the waveguide or cavity; meanwhile, the indirect coupling between meta-atoms via the feeds is captured by the second term on the right hand side in Eq. (8a).
@@
The commonly assumed simplified DMA model with a linear mapping from configuration to radiation pattern [27, 15, 28, 29, 19, 18, 30, 31, 20, 32] arises from assumptions of negligible mutual coupling and can be recovered from our physics-consistent model in the limiting case in which $\tilde{\mathbf{W}}$ vanishes:
@@
$\lim\limits_{\bar{\mathbf{W}}\to{\mathbf{0}_{N_{\mathrm{M}}}}}\!\!\!\left[e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})\right]\!=\!\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\!\!\operatorname{diag}(\mathbf{c}^{\prime})\mathbf{W}^{0}_{\mathcal{MF}}\!\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}\!\!\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}}$ (11)
@@
where the $i$th entry of $\mathbf{c}^{\prime}$ is the inverse of the $i$th entry of $\mathbf{c}$. In simplified models, the entries of $\mathbf{W}^{0}_{\mathcal{MF}}$ are identified based on simple 1D waveguide transmission properties, and $\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}$ is approximated as an identity matrix for stacked-microstrip DMAs because each feed is on a separate microstrip (see Fig. 1) and the microstrips are assumed to not couple to each other. An alternative route to recover the common simplified model from our physics-consistent model consists in imposing simultaneously that $G_{ij}=0\ \forall\ \{i,j\}\in\mathcal{M}$ and $G_{ij}\neq G_{ij}=0\ \forall\ i\in\mathcal{F},j\in\mathcal{M}$. The latter is the “unilateral approximation” [51] that prevents coupling between the meta-atoms via bounces off the feeds. (Note that this approximation introduces a non-reciprocity.) Clearly, by construction these simplified models cannot capture potential benefits of mutual coupling in optimizing DMAs for wireless communications.
@@
To systematically investigate the latter, a quantification of the DMA’s mutual coupling strength is required. Based on our system model insights, we quantify the DMA’s absolute mutual coupling strength with
@@
The ease of inverting $\tilde{\mathbf{W}}-\mathrm{diag}(\mathbf{c})$, and hence the strength of the non-linearity in the mapping from $\mathbf{c}$ to $\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})$ and ultimately $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$, depends on the DMA’s relative mutual coupling strength:
@@ II-D Summary
In this section, we have introduced a compact and implementation-agnostic DMA system model that is consistent with physics. We investigated how the model’s mathematical structure results in intertwined influences of the meta-atoms on the radiation pattern. We also clarified how to recover common simplified models from our physics-consistent model by neglecting mutual coupling. Finally, we introduced a metric to quantify the mutual coupling strength.
@@ III Theory
Equipped with the system model from Sec. II, in this section we theoretically analyze the strength and simplicity of the dependence of the DMA radiation pattern $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ on the DMA configuration $\mathbf{c}$. To that end, we analytically derive the first and second partial derivatives of $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ with respect to $\mathbf{c}$. The first derivative (Jacobian) informs us about the sensitivity of $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ to $\mathbf{c}$. The larger the Jacobian is, the stronger $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ depends on $\mathbf{c}$. We hypothesize that a larger sensitivity translates into a larger ability to tailor $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ to a desired wireless functionality; we confirm this hypothesis numerically and experimentally in the subsequent sections. The second derivative (Hessian) informs us about how the dependence of $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ on a given entry of $\mathbf{c}$ depends itself on another given entry of $\mathbf{c}$. The Hessian is the first term in the Taylor expansion of $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ with respect to $\mathbf{c}$ that indicates curvature. Thus, the larger the Hessian is, the less simple (i.e., more non-linear) the mapping from $\mathbf{c}$ to $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ is. For notational ease, we consider a DMA with a single feed (i.e., $N_{\mathrm{F}}=1$) in this section, but the underlying physical principles extend to DMAs with multiple feeds. For $N_{\mathrm{F}}=1$, $\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}}$ is a scalar that we denote by $e^{\mathrm{inc}}$.
@@ III-A Sensitivity of Radiation Pattern to Configuration
$\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})=\frac{\partial e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})}{\partial\mathbf{c}}=\mathbf{v}_{k}^{\top}(\mathbf{r})\mathbf{J}^{\mathrm{p}}(\mathbf{c})e^{\mathrm{inc}}\in\mathbb{C}^{N_{\mathrm{M}}},$ (14)
where $\mathbf{J}^{\mathrm{p}}(\mathbf{c})\in\mathbb{C}^{N_{\mathrm{M}}\times N_{\mathrm{M}}}$ is the Jacobian of the meta-atoms’ dipole moment, i.e., the $i$th column of $\mathbf{J}^{\mathrm{p}}(\mathbf{c})$ is $\mathbf{j}^{\mathrm{p}}_{i}(\mathbf{c})=\frac{\partial}{\partial c_{i}}\left(\left[\left(\mathbf{W}(\mathbf{c})\right)^{-1}\right]_{\mathcal{MF}}\right)\in\mathbb{C}^{N_{\mathrm{M}}}$ (recall that we chose $N_{\mathrm{F}}=1$). For notational ease, we define $\mathbf{\Omega}(\mathbf{c})=\left(\mathbf{W}(\mathbf{c})\right)^{-1}$. After a few linear algebra steps detailed in Appendix A, we obtain
$\mathbf{j}^{\mathrm{p}}_{i}(\mathbf{c})=\mathbf{\Omega}_{\mathcal{M}\mathcal{M}_{i}}(\mathbf{c})\mathbf{\Omega}_{\mathcal{M}_{i}\mathcal{F}}(\mathbf{c})=\mathbf{\Omega}_{\mathcal{M}\mathcal{M}_{i}}(\mathbf{c})\mathbf{p}_{\mathcal{M}_{i}}(\mathbf{c}),$ (15a)
$\mathbf{J}^{\mathrm{p}}(\mathbf{c})=\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})\mathrm{diag}(\mathbf{\Omega}_{\mathcal{MF}}(\mathbf{c}))=\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})\mathrm{diag}(\mathbf{p}_{\mathcal{M}}(\mathbf{c})).$ (15b)
For clarity, we recall that in Eq. (15) the cardinality of the sets $\mathcal{M}_{i}$ and $\mathcal{F}$ is unity, and that $\mathbf{\Omega}_{\mathcal{MF}}(\mathbf{c})=\mathbf{p}_{\mathcal{M}}(\mathbf{c})$. Substituting Eq. (15b) into Eq. (14), we have
$\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})=\mathbf{v}_{k}^{\top}(\mathbf{r})\mathbf{\Omega}_{\mathcal{M}\mathcal{M}}(\mathbf{c})\mathrm{diag}(\mathbf{p}_{\mathcal{M}}(\mathbf{c}))e^{\mathrm{inc}},$ (16)
@@
The absolute radiation pattern sensitivity can hence be quantified by $\|\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})\|_{2}$ based on Eq. (16). However, at this stage, some care is required to ensure a fair comparison between DMAs with different mutual coupling strengths because they likely differ regarding their radiation efficiency. Our sensitivity metric should be agnostic to the DMA’s radiation efficiency so that it informs us about the relative (as opposed to absolute) change of the radiation pattern in response to a change in DMA configuration. Thus, a normalization of $\|\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})\|_{2}$ is necessary.
@@
By definition, $\tilde{\mu}>0$; moreover, according to our numerical results in the next section, $\tilde{\mu}$ remains below unity in typical DMA architectures that we consider. For $10<\tilde{\mu}<$ 1, the larger $\tilde{\mu}$ is, the higher is the upper bound on $\|\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})\|_{2}$ according to Eq. (18) and hence on $\|\hat{\mathbf{J}}^{\mathrm{p}}(\mathbf{c})\|_{2}$ according to Eq. (17). In other words, the upper bound on the sensitivity increases with the relative mutual coupling strength. We further note that the upper bound is proportional to $\sigma_{N_{\mathrm{M}}}$ and thus to the largest “dressed” polarizability (which combines the effect of the polarizability and the self-interactions originating from $G_{ii}$ [52]) of the meta-atoms.
As explained in Appendix B, the bound could be attained with equality if the coupling between the meta-atoms was controllable. One possible route toward such controllable couplings would be the realization of a “beyond-diagonal DMA” (similar to BD-RIS [38, 53]) equipped with additional tunable lumped elements inside the cavity that do not radiate to the scene. We will explore BD-DMAs in future work.
Altogether, the theoretical analysis in this subsection suggests that the dependence of the DMA radiation pattern on the DMA configuration gets stronger as mutual coupling in the DMA gets stronger, even after normalization to remove the effect of the DMA’s radiation efficiency. We confirm this relation through extensive numerical and experimental studies in the next sections.

### v3

@@ II-C Insights into MC
$\begin{split}\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})&=\left[\left(\mathbf{W}(\mathbf{c})\right)^{-1}\right]_{\mathcal{MM}}=\left(\tilde{\mathbf{W}}-\mathrm{diag}(\mathbf{c})\right)^{-1}\\ &=\Bigl[\tilde{\mathbf{W}}^{-1}\!+\!\tilde{\mathbf{W}}^{-1}\bigl([\mathrm{diag}(\mathbf{c})]^{-1}\!-\!\tilde{\mathbf{W}}^{-1}\bigr)^{-1}\tilde{\mathbf{W}}^{-1}\Bigr],\end{split}$ (8b)
@@
$\bigl[\mathbf{W}(\mathbf{c})^{-1}\bigr]_{\mathcal{MF}}=-\,\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})\,\mathbf{W}^{0}_{\mathcal{MF}}\,\bigl(\mathbf{W}^{0}_{\mathcal{FF}}\bigr)^{-1}.$ (9)
Substituting Eq. (9) into Eq. (7) and using the expression in the second line of Eq. (8b) for $\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})$, we obtain
@@
$\begin{split}e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})=-\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\!\tilde{\mathbf{W}}^{-1}\mathbf{W}^{0}_{\mathcal{MF}}\!\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}\!\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}}\\ +\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\!\tilde{\mathbf{W}}^{-1}\!\bigl(\!\tilde{\mathbf{W}}^{-1}-\!\mathrm{diag}(\mathbf{c})^{-1}\bigr)^{-1}\tilde{\mathbf{W}}^{-1}\mathbf{W}^{0}_{\mathcal{MF}}\!\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}\!\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}}.\end{split}$ (10)
@@
The first [second] term on the right hand side in Eq. (10) represents contributions to the radiated field that are independent [dependent] of the DMA configuration $\mathbf{c}$. The dependence on $\mathbf{c}$ is seen to generally be non-linear due to a matrix inversion. By inspection of Eq. (8b), we observe that the off-diagonal elements of $\tilde{\mathbf{W}}$ mediate the coupling between the different entries of $\mathbf{c}$. This mathematical observation precisely matches the physical interpretation. The off-diagonal elements of the first term on the right hand side in Eq. (8a) contain the background Green’s functions between the meta-atoms which capture the coupling between them that is mediated via the waveguide or cavity; meanwhile, the indirect coupling between meta-atoms via the feeds is captured by the second term on the right hand side in Eq. (8a).
@@
The commonly assumed simplified DMA model with a linear mapping from configuration to radiation pattern [27, 14, 28, 29, 18, 17, 30, 31, 19, 32] arises from assumptions of negligible MC and can be recovered from our physics-consistent model in the limiting case in which the off-diagonal entries of $\tilde{\mathbf{W}}$ vanish. The first line of Eq. (8b) reveals that
@@
$\lim\limits_{\tilde{\mathbf{W}}\to{\mathrm{diag}(\mathbf{d})}}\!\!\!\!\!\!\left[\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})\right]=-\left[\mathrm{diag}(\mathbf{c}-\mathbf{d})\right]^{-1}=-\mathrm{diag}(\mathbf{c}^{\prime}),$ (11)
where $\mathbf{d}\in\mathbb{C}^{N_{\mathrm{M}}}$ is a vector containing the diagonal entries of $\tilde{\mathbf{W}}$, and the $i$th entry of $\mathbf{c}^{\prime}$ is the inverse of the difference between the $i$th entry of $\mathbf{c}$ and the $i$th entry of $\mathbf{d}$. Hence, in the limit of vanishing MC we recover the common simplified DMA model:
$\lim\limits_{\tilde{\mathbf{W}}\to{\mathrm{diag}(\mathbf{d})}}\!\!\!\left[e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})\right]\!=\!\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\!\!\mathrm{diag}(\mathbf{c}^{\prime})\mathbf{W}^{0}_{\mathcal{MF}}\!\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}\!\!\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}}.$ (12)
@@
In applications of Eq. (12) to stacked-microstrip DMAs, the entries of $\mathbf{W}^{0}_{\mathcal{MF}}$ are usually identified based on simple 1D waveguide transmission properties, and $\left(\mathbf{W}^{0}_{\mathcal{FF}}\right)^{-1}$ is approximated as an identity matrix because each feed is on a separate microstrip (see Fig. 1) and the microstrips are assumed not to couple to each other. An alternative route to recover the common simplified model from our physics-consistent model consists in imposing simultaneously that $G_{ij}=0$ for $i,j\in\mathcal{M},\ i\neq j$ and $G_{ij}=0$ for $i\in\mathcal{F},\ j\in\mathcal{M}$ (while $G_{ji}\neq 0$ for $j\in\mathcal{M},\ i\in\mathcal{F}$). This alternative route resembles the “unilateral approximation” [54] that prevents coupling between the meta-atoms via bounces off the feeds. (Note that this approximation introduces a non-reciprocity.) Clearly, by construction these simplified models cannot capture potential benefits of MC in optimizing DMAs for wireless communications.
@@
To systematically investigate the latter, a quantification of the DMA’s MC strength is required. Based on our system model insights, we quantify the DMA’s absolute MC strength with
@@
The ease of inverting $\tilde{\mathbf{W}}-\mathrm{diag}(\mathbf{c})$, and hence the strength of the non-linearity in the mapping from $\mathbf{c}$ to $\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})$ and ultimately $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$, depends on the DMA’s relative MC strength:
@@ II-D Summary
In this section, we have introduced a compact and implementation-agnostic DMA system model that is consistent with physics. We investigated how the model’s mathematical structure results in intertwined influences of the meta-atoms on the radiation pattern. We also clarified how to recover common simplified models from our physics-consistent model by neglecting MC. Finally, we introduced a metric to quantify the MC strength.
@@ III Theory
Equipped with the system model from Sec. II, in this section we theoretically analyze the strength and simplicity of the dependence of the DMA radiation pattern $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ on the DMA configuration $\mathbf{c}$. To that end, we analytically derive the first and second partial derivatives of $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ with respect to $\mathbf{c}$. The first derivative (Jacobian) informs us about the sensitivity of $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ to $\mathbf{c}$. The larger the Jacobian is, the stronger $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ depends on $\mathbf{c}$. We hypothesize that a larger sensitivity translates into a larger ability to tailor $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ to a desired wireless functionality; we confirm this hypothesis numerically and experimentally in the subsequent sections. The second derivative (Hessian) informs us about how the dependence of $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ on a given entry of $\mathbf{c}$ depends itself on another given entry of $\mathbf{c}$. The Hessian is the first term in the Taylor expansion of $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ with respect to $\mathbf{c}$ that indicates curvature. Thus, the larger the Hessian is, the less simple (i.e., more non-linear) the mapping from $\mathbf{c}$ to $e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})$ is.
For simplicity, we consider from now on a DMA with a single feed (i.e., $N_{\mathrm{F}}=1$). The analysis of the multiple-feed case is deferred to future work due to the complexity of considering the influence of the precoder on the partial derivative of the radiation pattern with respect to the DMA configuration. We anticipate that the key physical insights presented here will carry over to the multi-feed case. For $N_{\mathrm{F}}=1$, $\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}}$ is a scalar that we denote by $e^{\mathrm{inc}}$.
@@ III-A Sensitivity of Radiation Pattern to Configuration
$\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})=\frac{\partial e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})}{\partial\mathbf{c}}=\mathbf{v}_{k}^{\top}(\mathbf{r})\mathbf{J}^{\mathrm{p}}(\mathbf{c})\in\mathbb{C}^{N_{\mathrm{M}}},$ (15)
where $\mathbf{J}^{\mathrm{p}}(\mathbf{c})\in\mathbb{C}^{N_{\mathrm{M}}\times N_{\mathrm{M}}}$ is the Jacobian of the meta-atoms’ dipole moment, i.e., the $i$th column of $\mathbf{J}^{\mathrm{p}}(\mathbf{c})$ is $\mathbf{j}^{\mathrm{p}}_{i}(\mathbf{c})=\frac{\partial}{\partial c_{i}}\left(\left[\left(\mathbf{W}(\mathbf{c})\right)^{-1}\right]_{\mathcal{MF}}\right)e^{\mathrm{inc}}\in\mathbb{C}^{N_{\mathrm{M}}}$ (recall that we chose $N_{\mathrm{F}}=1$). For notational ease, we define $\mathbf{\Omega}(\mathbf{c})=\left(\mathbf{W}(\mathbf{c})\right)^{-1}$. After a few linear algebra steps detailed in Appendix A, we obtain
$\mathbf{j}^{\mathrm{p}}_{i}(\mathbf{c})=\mathbf{\Omega}_{\mathcal{M}\mathcal{M}_{i}}(\mathbf{c})\mathbf{p}_{\mathcal{M}_{i}}(\mathbf{c}),$ (16a)
$\mathbf{J}^{\mathrm{p}}(\mathbf{c})=\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})\mathrm{diag}(\mathbf{p}_{\mathcal{M}}(\mathbf{c})).$ (16b)
For clarity, we recall that in Eq. (16) the cardinality of the sets $\mathcal{M}_{i}$ and $\mathcal{F}$ is unity, and that $\mathbf{p}_{\mathcal{M}}(\mathbf{c})=\mathbf{\Omega}_{\mathcal{MF}}(\mathbf{c})e^{\mathrm{inc}}$. Substituting Eq. (16b) into Eq. (15), we have
$\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})=\mathbf{v}_{k}^{\top}(\mathbf{r})\mathbf{\Omega}_{\mathcal{M}\mathcal{M}}(\mathbf{c})\mathrm{diag}(\mathbf{p}_{\mathcal{M}}(\mathbf{c})),$ (17)
@@
The absolute radiation pattern sensitivity can hence be quantified by $\|\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})\|_{2}$ based on Eq. (17). However, at this stage, some care is required to ensure a fair comparison between DMAs with different MC strengths because they likely differ regarding their radiation efficiency. Our sensitivity metric should be agnostic to the DMA’s radiation efficiency so that it informs us about the relative (as opposed to absolute) change of the radiation pattern in response to a change in DMA configuration. Thus, a normalization of $\|\mathbf{j}^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})\|_{2}$ is necessary.
@@
By definition, $\tilde{\mu}>0$; moreover, according to our numerical results in the next section, $\tilde{\mu}$ remains below unity in typical DMA architectures that we consider. For $0<\tilde{\mu}<1$, the larger $\tilde{\mu}$ is, the higher is the upper bound on $\|\mathbf{\Omega}_{\mathcal{MM}}(\mathbf{c})\|_{2}$ according to Eq. (19) and hence on $\|\hat{\mathbf{J}}^{\mathrm{p}}(\mathbf{c})\|_{2}$ according to Eq. (18). In other words, the upper bound on the sensitivity increases with the relative MC strength. We further note that the upper bound is inversely proportional to $\sigma_{N_{\mathrm{M}}}$, and thus it increases with the largest “dressed” polarizability (which combines the effect of the polarizability and the self-interactions originating from $G_{ii}$ [55]) of the meta-atoms.
As explained in Appendix B, the bound could be attained with equality if the coupling between the meta-atoms was controllable. One possible route toward such controllable couplings would be the realization of a “beyond-diagonal DMA” (similar to BD-RIS [38, 56]) equipped with additional tunable lumped elements inside the cavity that do not radiate to the scene. An initial exploration of BD-DMAs can be found in our recent follow-up work [57].
Altogether, the theoretical analysis in this subsection suggests that the dependence of the DMA radiation pattern on the DMA configuration gets stronger as MC in the DMA gets stronger, even after normalization to remove the effect of the DMA’s radiation efficiency. We confirm this relation through extensive numerical and experimental studies in the next sections.